OpenStudy (dumbsearch2):

If 90n + 23p = 4523, and n and p are both positive integers, what is one possible value of n + p?

OpenStudy (anonymous):

n = 50, p = 1

OpenStudy (anonymous):

(27,91) I just graph p = 4523/23 -( 90n/23)

OpenStudy (anonymous):

oh n+ p = 51 or n+p = 118

OpenStudy (anonymous):

it's funny how I forgot to answer the initial question xD

OpenStudy (anonymous):

you will have to test p=1,11,21,31,41,51,61,71,81,91,101,111, 121,... because 4523-23p=90n=3.3.10.n it's a ten and nine multiple

OpenStudy (anonymous):

p=(1,91,181) and n=(50,27,4)

OpenStudy (anonymous):

@dumbsearch2 get it?